Find an equation of the line that contains the point(4,-9) and is parallel to the line that contains(-4,8) and (8,-1) enter your answer in the form y=mx+b unfactored do not use decimals leave improper fractions improper
Answer by Alan3354(69443)

Same problem, different numbers.
---------------------------------------------------
Find the equation of the line that contains the point P(-4,18) and is parallel to the line that contains A(2,-2) and B(6,3).
----------
The slope of AB, m = diffy/diffx
m = 5/4
-------------
Parallel lines have the same slope.
y-y1 = m*(x-x1) where (x1,y1) is the point P.
y - 18 = (5/4)*(x+4)
y = (5/4)x + 23